OpenStudy (anonymous):

how do i find the limit [(x^4)-1] / [(x^3)-1] as x approaches 1 ?

OpenStudy (chaise):

Substitute your x = 1 values into your limit, if you get an answer which is "0/0" you must factorise and cancel, if you get "x/b" then this is your answer, if you get "0/x" then there is no limit.

OpenStudy (anonymous):

x^4-1=(x-1)(x+1)(x^2+1); x^3-1=(x-1)(x^2+x+1); So limit is 4/3

OpenStudy (chaise):

This is right.

OpenStudy (anonymous):

great..thanks

OpenStudy (anonymous):

Not at all.